0 0
Add a comment Improve this question Transcribed image text
Answer #1

A conditional jump is called a branch in MIPS. The offset is a 16-bit value. It can be computed as follows:

Offset = target instruction - (branch instruction + 1).

Now let us number each instruction starting form 0 as follows.

0 loop: beq $s0, $s1, done

  1. ….
  2. done: beq $t0, $t1, skip
  3. skip: addi ….
  4. beq $t1, $t5, done
  5. jend

offset of instruction 0 = 5 – (0 + 1) = 4

offset of instruction 5 = 8 – (5 + 1) = 2

offset of instruction 9 = 5 – (9 + 1) = 5